Camile owns Crunch Code, a company that provides quick programming solutions. Clients send Crunch projects via their webpage and Crunch bangs out the needed code as quickly as possible. Camile has five programmers who do all of the coding. On average, a project arrives once every 4.8 hours. Each project is assigned to one programmer and that programmer takes on average 19.2 hours to complete each project.

Camile wants to know what fraction (as a percent) of the time is each of her programmers busy coding?
Answer #1

Avg Utilzation= Arrival rate/(servers*service rate)

Arrival rate= 1/4.8=0.2083 per hr

service rate=1/19.2=0.0521 per hour

hence avg utilizarion of coders= 0.2083/(5*0.0521)=0.799=0.8 or 80%